# LED: Light Enhanced Depth Estimation at Night

### v2

@@ Abstract
Nighttime camera-based depth estimation is a highly challenging task, especially for autonomous driving applications, where accurate depth perception is essential for ensuring safe navigation. We aim to improve the reliability of perception systems at night time, where models trained on daytime data often fail in the absence of precise but costly LiDAR sensors. In this work, we introduce Light Enhanced Depth (LED), a novel cost-effective approach that significantly improves depth estimation in low-light environments by harnessing a pattern projected by high definition headlights available in modern vehicles. LED leads to significant performance boosts across multiple depth-estimation architectures (encoder-decoder, Adabins, DepthFormer) both on synthetic and real datasets. Furthermore, increased performances beyond illuminated areas reveal a holistic enhancement in scene understanding. Finally, we release the Nighttime Synthetic Drive Dataset, a new synthetic and photo-realistic nighttime dataset, which comprises 49,990 comprehensively annotated images. To facilitate further research, both synthetic dataset and code are publicly available at https://simondemoreau.github.io/LED/.
11footnotetext: Center for Robotics - Mines Paris - PSL University, France, {simon.de_moreau, bodgan.stanciulescu, fabien.moutarde} @minesparis.psl.eu22footnotetext: Valeo, France, {simon.de-moreau, yasser.almehio, hafid.el-idrissi} @valeo.com33footnotetext: Valeo.ai, France, andrei.bursuc@valeo.com
@@ 1 Introduction
Adverse conditions, such as harsh weather or nighttime, pose significant challenges to many computer vision applications. Despite impressive progress in perception systems for autonomous driving, enabled by powerful deep neural architectures and training techniques, the challenges of nighttime navigation persist, demanding precise scene comprehension. Accurate depth estimation is a crucial aspect of perception and extends beyond immediate applications, profoundly impacting overall scene perception, especially at night [37]. While LiDAR sensors offer high accuracy, their widespread adoption is impeded by substantial costs. Additionally, in humid weather such as fog or rain, LiDAR effectiveness can be significantly degraded due to beam reflection and increased noise levels [6, 7].
Cameras are powerful and cost-effective sensors capturing rich information about the environment while deployed in most modern-day vehicles. Camera-based perception solutions displays high reliability and accuracy [16, 5, 17, 15, 4, 1, 23], but mostly on clear daylight conditions. However, they struggle under distribution shift and low-light conditions. An additional challenge is related to the availability of labeled data for training high-capacity deep neural networks. For the task of depth estimation, some methods rely on supervised learning [4, 1, 23, 20], most use self-supervised methods [16, 32, 48, 46, 36, 5, 17, 45, 47] since obtaining dense ground truth depth information can be expensive. While dedicated approaches for particular conditions arised in the past few years [45, 47, 40, 38, 24] they are still partially mitigating these issues. To the best of our knowledge, they remain limited to self-supervised strategies due to the absence of publicly available large-scale annotated datasets suitable for nighttime depth estimation training. This paper introduces a new annotated nighttime synthetic dataset featuring dense depth maps, among other labels.
@@
Figure 1: LED (Light Enhanced Depth) is a novel method enhancing depth estimation from nighttime RGB images by harnessing High-Definition headlights’ pattern. We release a new synthetic nighttime dataset containing both high beam and pattern-illuminated RGB images, associated with comprehensive ground truth annotations. This dataset will facilitate research into nighttime perception across various applications.
To improve the performance of supervised methods in low-light conditions, we draw inspiration from active stereovision methods [13, 33, 21, 2] by harnessing High-Definition (HD) headlights commonly equipped in modern vehicles. These headlights have demonstrated promising results in scene perception research [42, 43, 41]. Therefore, we leverage them to project a pattern into the scene, guiding the network and thereby improving perception accuracy and reliability.
In this paper, we introduce Light Enhanced Depth (LED), a novel approach that significantly improves depth estimation in low-light environments, ensuring enhanced accuracy and robust performances for autonomous vehicles. Our contributions can be summarized as follows:
  • Enhanced Depth Estimation: Using only an encoder-decoder as depth-estimation model, we demonstrate a remarkable -11% RMSE improvement at nighttime by incorporating HD headlight pattern.
  • Architecture-Agnostic: We extend our methodology to two other state-of-the-art (SOTA) methods designed for daytime scenarios, showcasing its versatility and performance improvement across diverse neural network architectures (-24.06% RMSE on Adabins [4], -8.00% RMSE on DepthFormer [23]).
  • Global Scene Understanding: We show that our proposed method enhances the overall understanding of the scene beyond illuminated area with -3.05% RMSE on the whole image.
  • Dataset: We provide the Nighttime Synthetic Drive Dataset, a photorealistic synthetic dataset comprising 49,995 comprehensively annotated nighttime images to foster future investigations in nighttime perception.
@@
  • Real prototype: LED also demonstrates great performances improvements on our in-house dataset, collected using a real car-mounted prototype.
@@ 2 Related Work
2.1 Daytime Depth Estimation Methods at Night
Supervised Learning from Monocular Camera.
Recent advances in supervised depth prediction exploit transformer architectures and attention mechanisms [4, 1, 23, 20] to improve performances. These methods highlight the effectiveness of transformer-based architectures and the strategic use of bins [4, 1] in improving depth prediction accuracy, even with more limited datasets available [14, 9, 27, 29] compared to self-supervised methods [16, 32, 48, 46, 5, 17, 36].
However, to the best of our knowledge, these methods are confined to daytime scenarios. Nighttime depth estimation approaches [45, 47] currently rely on self-supervised methods. This limitation stems from the scarcity of large-scale nighttime datasets and, even more so, the lack of properly annotated ones. In response, we release the Nighttime Synthetic Drive Dataset, to support the development of approaches adapted to nighttime scenarios.
Nighttime Domain Adaptation.
[40] and [38] tackle nighttime depth estimation as a domain adaptation challenge, focusing on common feature representations for daytime and nighttime images. The former maintains performance by aligning night image encoder features with pre-trained daytime encoder features. Similarly, the latter enhances nighttime depth estimation by leveraging dense features representation across domains. In contrast, [24] reject the assumption of identical features in both domains. To capture their nuances, they train independent feature extractors with losses considering domains similarities and differences. With LED, we tailor our method to fit the unique features of nighttime scenes by leveraging the informative deformations of the light pattern emitted by a vehicle’s HD headlamp.
2.2 Specific Modalities for Nighttime
LiDAR or thermal imaging [26, 35] can be used to overcome low-light challenges. However, these approaches require paired data for training and entail extra expenses for sensors and computational power needed for vehicle integration. In this work, we address nighttime challenges by leveraging built-in equipment available in modern vehicles.
@@ 2.3 Light for Perception
Active Stereovision.
Depth estimation based on active stereovision involves disparity measurement. Unlike traditional stereovision, disparity is computed between an image and a pattern projected onto the scene. Recent methods [13, 33, 21, 2] use deep learning models designed to take patterns and images as input. They propose specific loss functions dedicated to disparity maps.
While active stereovision offers valuable scale information, it has demonstrated significant performance degradation when deployed outdoors, due to high ambient lighting and the low power of projectors [28, 18, 8]. We avoid this degradation thanks to our nighttime environment characterized by low illumination conditions and the utilization of high-powered visible HD headlights, resulting in favorable conditions for our approach.
HD Lighting.
HD lighting has been a research interest for years, but its recent industrialization explains the limited number of related works. A series of articles [11, 39] highlight its potential applications and particularly its ability to design anti-glare systems while optimizing illumination for enhanced driver visibility. They have also demonstrated improved driver visibility in adverse conditions such as rain or snow. More recently, [44] proposed reducing overall scene illumination to decrease power consumption while maintaining object detection capabilities. To facilitate these experiments, the same team has proposed hardware-in-the-loop simulation solutions for headlights [42, 43, 41].
We introduce a novel application of HD headlights to improve understanding of overall scene geometry and enhance depth estimation. We achieve this by projecting a pattern onto the nighttime scene, providing guidance for the model.
Figure 2: Pattern deformation and size variation in the scene: left image shows the pattern projection on a depth image, computed using geometrical rules. It demonstrates trapezoidal deformations on horizontal surfaces and undistorted squares on vertical surfaces. Complex deformations on the car reveal insights into its geometry. The right image showcases the pattern size projected on a wall at 10 m (darker background) and 100 m (lighter background). Note that square sizes increase with distance.
@@ 3.1 HD Pattern and Headlights
Networks designed for active stereovision exploit the disparity between the projected pattern and the camera’s view to estimate distances within the scene. In LED, a model identifies pattern’s areas that deviate from the implicitly learned reference pattern.
Diverging from traditional active stereovision which relies on infrared lasers projecting points, we harness an HD headlight. Similar to a beamer, this type of headlight can dynamically project any image or shape onto the scene. Specifically, we simulate a real 3,696 pixels HD headlight (132$\times$28 px) with a horizontal Field of View (FOV) of 35° and a vertical FOV of 7°. The emitted light is not colimated, thus, similar to a projector, the size of a pixel in the projection varies with distance. We illustrate it in Fig. 2, as the projection extends further, a pixel appears larger. This provides the neural network with valuable cues about distances in the scene. To avoid superposition of HD patterns we only use the left headlight.
We employ a regular checkerboard pattern because of its ability to exhibit high discontinuities and strong contrast. The dense concentration of corners and transitions in the pattern serves as features easily recognizable and detectable by convolutional neural networks [3]. These features aid the network in extracting the pattern from the scene. Dimensions of the checkerboard cells are 0.5°, ensuring proper visibility with the simulated camera resolution.
Upon projection, the checkerboard undergoes deformations (from the camera’s perspective) based on surface’s shape. As illustrated in Fig. 2, horizontal planes tend to stretch the pattern into a trapezoid, while vertical surfaces parallel to image plane do not cause any deformation. More complex deformations occur on non-planar surfaces such as cars. All these elements provide valuable information about the scene’s geometry to the network and thus enhances the quality of depth estimation.
@@ 3.2 Architectures
Our method is architecture-agnostic. Since we rely on a single pattern, the model can learn the reference pattern implicitly. Consequently, there is no need to design a dedicated network architecture to handle it.
To showcase the benefits of using a projected HD pattern for depth estimation, we employ an encoder-decoder model (U-Net [34]) for most of our studies. This model takes a single RGB image with the projected pattern as input and estimates depth map as output. Despite the simplicity of this architecture, it achieve results on par with or superior to other SOTA methods (see Sec. 5.1).
We successfully apply our method to more complex SOTA architectures: DepthFormer [23] and Adabins [4], with similar performances enhancement (see Sec. 5.4).

### v3

@@ Abstract
Nighttime camera-based depth estimation is a highly challenging task, especially for autonomous driving applications, where accurate depth perception is essential for ensuring safe navigation. Models trained on daytime data often fail in the absence of precise but costly LiDAR. Even vision foundation models trained on large amounts of data are unreliable in low-light conditions. In this work, we aim to improve the reliability of perception systems at night time. To this end, we introduce Light Enhanced Depth (LED), a novel, cost-effective approach that significantly improves depth estimation in low-light environments by harnessing a pattern projected by high definition headlights available in modern vehicles. LED leads to significant performance boosts across multiple depth-estimation architectures (encoder-decoder, Adabins, DepthFormer, Depth Anything V2) both on synthetic and real datasets. Furthermore, increased performances beyond illuminated areas reveal a holistic enhancement in scene understanding. Finally, we release the Nighttime Synthetic Drive Dataset, a synthetic and photo-realistic nighttime dataset, which comprises 49,990 comprehensively annotated images. To facilitate further research, both synthetic dataset and code are publicly available at https://simondemoreau.github.io/LED/.
@@ 1 Introduction
Figure 1: Light Enhanced Depth (LED) is a novel method that harnesses high-definition headlights’ pattern projected onto the scene to enhance nighttime depth estimation from RGB images. We release a synthetic nighttime dataset with high beam and pattern-illuminated images, along with comprehensive ground truth annotations, to advance research in nighttime perception.
Adverse conditions, such as harsh weather or nighttime, pose significant challenges to many computer vision applications. Despite impressive progress in perception systems for autonomous driving, enabled by powerful deep neural architectures and training techniques, the challenges of nighttime navigation persist. Accurate depth estimation is a crucial aspect of perception, profoundly impacting overall scene perception and comprehension, especially at night [Silberman and Fergus(2011)]. While LiDAR sensors offer high accuracy, their widespread adoption is impeded by their substantial cost. Additionally, in humid weather, LiDAR effectiveness can be significantly degraded due to beam reflection and increased noise levels [Bijelic et al.(2018a)Bijelic, Gruber, and Ritter, Bijelic et al.(2018b)Bijelic, Gruber, and Ritter].
Cameras are powerful, cost-effective sensors that capture rich information about the environment and are deployed in most modern vehicles. Camera-based perception solutions display high reliability and accuracy [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow, Bian et al.(2019)Bian, Li, Wang, Zhan, Shen, Cheng, and Reid, Guizilini et al.(2020)Guizilini, Ambrus, Pillai, Raventos, and Gaidon, Godard et al.(2017)Godard, Mac Aodha, and Brostow, Bhat et al.(2021)Bhat, Alhashim, and Wonka, Agarwal and Arora(2023), Li et al.(2023)Li, Chen, Liu, and Jiang], but mostly on clear daylight conditions: they struggle under distribution shift and low-light conditions. While foundation models [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao, Bochkovskii et al.(2024)Bochkovskii, Delaunoy, Germain, Santos, Zhou, Richter, and Koltun, Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] show great improvement for depth estimation on multiple domains, they are still not robust to nighttime images that are long-tail in the training data distribution. The availability of labeled data necessary for training high-capacity deep neural networks is an additional challenge. For the task of depth estimation, some methods rely on supervised learning [Bhat et al.(2021)Bhat, Alhashim, and Wonka, Agarwal and Arora(2023), Li et al.(2023)Li, Chen, Liu, and Jiang, Lee et al.(2019)Lee, Han, Ko, and Suh], but most use self-supervised methods [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow, Poggi et al.(2018)Poggi, Aleotti, Tosi, and Mattoccia, Zhou et al.(2017)Zhou, Brown, Snavely, and Lowe, Yin and Shi(2018), Shu et al.(2020)Shu, Yu, Duan, and Yang, Bian et al.(2019)Bian, Li, Wang, Zhan, Shen, Cheng, and Reid, Guizilini et al.(2020)Guizilini, Ambrus, Pillai, Raventos, and Gaidon, Wang et al.(2021)Wang, Zhang, Yan, Li, Xu, Li, and Yang, Zheng et al.(2023)Zheng, Zhong, Li, Gao, Zheng, Jin, Wang, Zhao, Zhou, Zhang, et al.] since obtaining ground truth depth information is expensive. While dedicated approaches for specific conditions have emerged in recent years [Wang et al.(2021)Wang, Zhang, Yan, Li, Xu, Li, and Yang, Zheng et al.(2023)Zheng, Zhong, Li, Gao, Zheng, Jin, Wang, Zhao, Zhou, Zhang, et al., Vankadari et al.(2020)Vankadari, Garg, Majumder, Kumar, and Behera, Spencer et al.(2020)Spencer, Bowden, and Hadfield, Liu et al.(2021)Liu, Song, Wang, Liu, and Zhang], they still only partially mitigate these issues. To the best of our knowledge, they remain limited to self-supervised strategies, due to the lack of publicly available, large-scale annotated datasets suitable for nighttime depth estimation. To address this gap, this paper releases a synthetic nighttime dataset annotated with dense depth maps, along with additional labels.
In this paper, we introduce Light Enhanced Depth (LED), a novel approach that significantly improves depth estimation in low-light environments, ensuring enhanced accuracy and reliability for autonomous vehicles. High-Definition (HD) headlights, commonly found in modern vehicles, have shown promising results in scene perception research [Waldner et al.(2019)Waldner, Kramer, and Bertram, Waldner et al.(2020)Waldner, Kramer, and Bertram, Waldner and Bertram(2021)]. Drawing inspiration from active stereovision [Fanello et al.(2016)Fanello, Rhemann, Tankovich, Kowdle, Escolano, Kim, and Izadi, Riegler et al.(2019)Riegler, Liao, Donne, Koltun, and Geiger, Li et al.(2022)Li, Monno, and Okutomi, Baek and Heide(2021)], we harness HD headlights to project a pattern into the scene, guiding the network and thereby improving performance.
Our contributions can be summarized as follows:
$\bullet$ Architecture-agnostic enhancement: LED can be applied to any depth estimation architecture to improve nighttime performances (RMSE: -11% on encoder-decoder, -24.06% on Adabins [Bhat et al.(2021)Bhat, Alhashim, and Wonka], -8.00% on DepthFormer [Li et al.(2023)Li, Chen, Liu, and Jiang], -15.5% on Depth Anything V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao]).
$\bullet$ Data-efficiency: LED-trained models outperform the others with only 20% of training data.
@@
$\bullet$ Real prototype: LED demonstrates great performances improvements on our in-house dataset, collected using a real car-mounted prototype.
$\bullet$ Dataset: We provide the Nighttime Synthetic Drive Dataset, a photorealistic synthetic dataset, comprising 49,995 comprehensively annotated nighttime images to foster future investigations in nighttime perception.
@@ 2 Related Work
2.1 Depth Estimation from Camera
Supervised Learning. Recent advances in supervised depth prediction leverage transformer architectures and attention mechanisms [Bhat et al.(2021)Bhat, Alhashim, and Wonka, Agarwal and Arora(2023), Li et al.(2023)Li, Chen, Liu, and Jiang, Lee et al.(2019)Lee, Han, Ko, and Suh] to improve performances. The use of bins [Bhat et al.(2021)Bhat, Alhashim, and Wonka, Agarwal and Arora(2023)] improves depth prediction accuracy even with limited datasets available [Geiger et al.(2012)Geiger, Lenz, and Urtasun, Caesar et al.(2020)Caesar, Bankiti, Lang, Vora, Liong, Xu, Krishnan, Pan, Baldan, and Beijbom, Maddern et al.(2017)Maddern, Pascoe, Linegar, and Newman, Nathan Silberman and Fergus(2012)] compared to self-supervised methods [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow, Poggi et al.(2018)Poggi, Aleotti, Tosi, and Mattoccia, Zhou et al.(2017)Zhou, Brown, Snavely, and Lowe, Yin and Shi(2018), Bian et al.(2019)Bian, Li, Wang, Zhan, Shen, Cheng, and Reid, Guizilini et al.(2020)Guizilini, Ambrus, Pillai, Raventos, and Gaidon, Shu et al.(2020)Shu, Yu, Duan, and Yang]. However, these methods are confined to daytime scenarios. Nighttime depth estimation approaches [Wang et al.(2021)Wang, Zhang, Yan, Li, Xu, Li, and Yang, Zheng et al.(2023)Zheng, Zhong, Li, Gao, Zheng, Jin, Wang, Zhao, Zhou, Zhang, et al.] rely on self-supervised methods, due to the lack of large, annotated nighttime datasets. In response, we release the Nighttime Synthetic Drive Dataset to support the development of nighttime-specific methods.
Vision Foundation Models. Depth foundation models [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao, Bochkovskii et al.(2024)Bochkovskii, Delaunoy, Germain, Santos, Zhou, Richter, and Koltun, Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] have achieved impressive generalization by leveraging extensive synthetic and real-world training data. However, we show in section˜5.4 that these models perform poorly in zero-shot evaluations on our nighttime dataset. It highlights the need for models specialized in nighttime conditions and underscores the benefits of our method.
Nighttime Domain Adaptation. Nighttime depth estimation can be approached as a domain adaptation problem [Gasperini et al.(2023)Gasperini, Morbitzer, Jung, Navab, and Tombari], focusing on aligning features between daytime and nighttime images [Vankadari et al.(2020)Vankadari, Garg, Majumder, Kumar, and Behera, Spencer et al.(2020)Spencer, Bowden, and Hadfield, Liu et al.(2021)Liu, Song, Wang, Liu, and Zhang, Liu et al.(2021)Liu, Song, Wang, Liu, and Zhang].  Our method specifically adapts to nighttime scenes features by leveraging the informative deformations of the light pattern emitted by a vehicle’s HD headlamp.
@@ 2.2 Light for Perception
Active Stereovision. Depth estimation based on active stereovision involves disparity measurement. Unlike traditional stereovision, disparity is computed between an image and a pattern projected onto the scene. Recent methods [Fanello et al.(2016)Fanello, Rhemann, Tankovich, Kowdle, Escolano, Kim, and Izadi, Riegler et al.(2019)Riegler, Liao, Donne, Koltun, and Geiger, Li et al.(2022)Li, Monno, and Okutomi, Baek and Heide(2021)] use deep learning models designed to take patterns and images as input. While active stereovision offers valuable scale information, it has demonstrated significant performance degradation when deployed outdoors, due to high ambient lighting and the low power of projectors [Mertz et al.(2012)Mertz, Koppal, Sia, and Narasimhan, Gupta et al.(2013)Gupta, Yin, and Nayar, Brahmanage and Leung(2019)]. We avoid this thanks to our nighttime environment and the utilization of high-powered HD headlights.
HD Lighting. A series of works [De Charette et al.(2012)De Charette, Tamburo, Barnum, Rowe, Kanade, and Narasimhan, Tamburo et al.(2014)Tamburo, Nurvitadhi, Chugh, Chen, Rowe, Kanade, and Narasimhan] highlight HD lighting potential applications and particularly its ability to design anti-glare systems while optimizing illumination for enhanced driver visibility, even in adverse conditions such as rain or snow. More recently, [Waldner et al.(2022)Waldner, Müller, and Bertram] propose reducing overall scene illumination to decrease power consumption while maintaining object detection capabilities. [Waldner et al.(2019)Waldner, Kramer, and Bertram, Waldner et al.(2020)Waldner, Kramer, and Bertram, Waldner and Bertram(2021)] also have proposed hardware-in-the-loop simulation for HD headlights. We introduce a novel application of HD headlights to improve understanding of overall scene geometry and enhance depth estimation. We achieve this by projecting a pattern onto the nighttime scene, providing guidance for the model.
@@ 3.1 HD Pattern and Headlights
Networks designed for active stereovision exploit the disparity between the projected pattern and the camera’s view to estimate distances within the scene. With LED, a model identifies pattern’s areas that deviate from the implicitly learned reference pattern.
Figure 2: Pattern deformation in the scene: Left: For explanatory purpose, we project the pattern on a depth image. It demonstrates trapezoidal deformations on horizontal surfaces and undistorted squares on vertical surfaces. Complex deformations on the car reveal insights into its geometry. Right: Pattern projected on a wall at 10 m (darker) and 100 m (lighter). Square sizes increases with distance.
Unlike traditional active stereovision, which relies on infrared lasers to project points, we harness an HD headlight. Similar to a projector, this headlight can dynamically project any image or shape onto the scene. To prevent pattern overlap, we use only the left headlight.
We employ a regular checkerboard pattern, because of its high contrast and sharp discontinuities. The dense concentration of corners and transitions makes it easily detectable by convolutional neural networks [Baker et al.(2020)Baker, Lu, Erlikhman, and Kellman]. Upon projection, the checkerboard is deformed according to surface shapes (from the camera’s perspective). Horizontal planes stretch the pattern into a trapezoid, while vertical surfaces parallel to the image plane cause no deformation (see fig.˜2). Non-planar surfaces, such as cars, result in more complex distortions. These deformations provide crucial geometric cues to the network, improving depth estimation accuracy. Additionally, the emitted light is not colimated: as the projection extends further, the light becomes more divergent, causing pixels to appear larger (see fig.˜2), thus providing useful depth cues to the network.
@@ 3.2 Architectures
Our method is architecture-agnostic. By relying on a single pattern, the model can implicitly learn it, eliminating the need for a dedicated network architecture. To demonstrate the benefits of projecting an HD pattern for depth estimation, we leverage a simple encoder-decoder model [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox]. In spite of its straight-forward architecture, it achieves results on par with or superior to other state-of-the-art (SOTA) methods (see section˜5.1). We also successfully apply our method to more complex SOTA architectures, such as DepthFormer [Li et al.(2023)Li, Chen, Liu, and Jiang] and Adabins [Bhat et al.(2021)Bhat, Alhashim, and Wonka], with similar performance gains (see section˜5.1). Finally, we show that finetuning Depth Anything V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] with LED, on a reduced amount of training samples, enable robust and precise depth estimation.
3.3 Implementation Details
We implement the encoder-decoder in PyTorch [Paszke et al.(2019)Paszke, Gross, Massa, Lerer, Bradbury, Chanan, Killeen, Lin, Gimelshein, Antiga, Desmaison, Kopf, Yang, DeVito, Raison, Tejani, Chilamkurthy, Steiner, Fang, Bai, and Chintala]. The model is trained from scratch during 70 epochs using the AdamW optimizer [Loshchilov and Hutter(2018)], with a batch size of 32 and a learning rate of $10^{-3}$. Input and output resolutions are set to $\times320$ px. Similar to [Hu et al.(2019)Hu, Ozay, Zhang, and Okatani], our learning objective combines losses based on Log L1, gradient and normals. More information is available in the supplementary. The selection of the best epoch is based on the Root Mean Square Error (RMSE). The training process takes 4 hours on a single Nvidia RTX 4090.
We implement Adabins [Bhat et al.(2021)Bhat, Alhashim, and Wonka] using the official code and rely on the toolbox [Li(2022)] for DepthFormer [Li et al.(2023)Li, Chen, Liu, and Jiang]. We apply Depth Anything V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] official fine-tuning code to adapt the model for metric depth estimation. The only changes are limited to dataset format compatibility. The training setup follows the original papers’ recommendations, excluding data augmentation to preserve the HD pattern from cropping.